. The debt ratio (debt/value) is.80. Total assets are $10 million. Find equity. Find the debt-equity ratio A firm has a debt/equity ratio of 3.00. Find the debt/value ratio. You can assume total assets $10 million.
Answer #1

a.Debt ratio=Debt/Total assets

Hence debt=(0.8*10)=$8 million

Total assets=debt+equity

Hence equity=(10-8)=$2million

Debt to equity ratio=Debt/equity

=(8/2)=4

b.Debt/equity ratio=Debt/equity

Hence debt=3equity

Total assets=Debt+Equity

Hence 10=3equity+equity

Hence equity=10/(3+1)=$2.5million

Hence debt=(3*2.5)=$7.5million

Hence debt/value ratio=(7.5/10)=0.75